OpenStudy (anonymous):

Can someone tell me if I got this problem correct and if not, what did I do wrong? For the pair of functions, f(x)= 3x and g(x) =2x^2-4x+5. Find the following: (a) ( f o g) (x) (b) (f o g) (-4) (c ) (g o f) (x) (d) (g o f) (-4) My answers: a. 6x^2-12x+15 b. 39 c. 18x^2-12x+5 d. -19

OpenStudy (raden):

a and c were right, but b and d still incorrect

OpenStudy (raden):

( f o g) (x) = 6x^2-12x+15 so, ( f o g) (-4) = 6(-4)^2 - 12(-4) + 15 = 6(16)+48+15 = 96+48+15=159

OpenStudy (raden):

(g o f) (x) = 18x^2-12x+5 (g o f) (-4) = 18(-4)^2-12(-4)+5 = 18(16)+48+5 = 288+48+5 = 341

OpenStudy (anonymous):

I see what I did wrong, thank you for your help

OpenStudy (raden):

yw